a) Find the potential at a point in space located at 0.07m from a charge -50 x10-6C. Find the potential energy if a charge +4 x10-6C is located at that point.



1
Expert's answer
2021-10-26T17:17:48-0400

"\varphi=\frac{kq_1}r,"

"\varphi=-\frac{9\cdot 10^9\cdot 50\cdot 10^{-6}}{0.07}=-6.4~MV,"

"W=-\varphi q_2,"

"W=6.4\cdot 10^6 \cdot 4\cdot 10^{-6}=25.6~J."